2 divided by 8x-6 is equal to 4 divided by 4x+2
Answer by checkley71(8403)

2/(8x-6)=4/(4x+2) or 4(8x-6)=2(4x+2) or 32x-24=8x+4 or 32x-8x=4+24 or
24x=28 or 6x=7 or x=7/6
proof 2/(8*7/6-6)=4/(4*7/6+2) or 2/(56/6-6)=4/(28/6+2) or
2/(56/6-36/6)=4/(28/6+12/6) or 4/(20/6)=2/(40/6) or 4(20/6)=2(40/6) or
80/6=80/6